OpenStudy (anonymous):

What is the solution of the system? Use substitution x=-2y+10 -6y=6x-6 z=-3x +10y

OpenStudy (zepp):

Using the first one x = -2y + 10 Substitute x into the 2nd equation -6y = 6 (-2y +10) -6 -6y = -12y + 60 - 6 -6y = -12y +54 -12y - -6y = 54 -12y + 6y = 54 -6y = 54 y = -9 Then x = -2y + 10 x = -2(-9) + 10 x = 18 + 10 x = 28 Can you go on?

OpenStudy (zepp):

z = -3x +10y z = -3(28) + 10(-9) z = -84 + -90 z = -84 - 90 z = -84-90 z = -174

OpenStudy (anonymous):

i have no idea what all the this is